The maximum kinetic energy of emitted photoelectrons in a photoelectric cell is 7.5X10-19 J. What is the stopping 

1
Expert's answer
2022-06-13T08:23:34-0400
"W=eV"

"V=\frac{W}{e}=\frac{7.5*10^{-19}\:\rm J}{1.6*10^{-19}\:\rm C}=4.7\:\rm V"